OpenStudy (anonymous):

B={7,8,9,10,11,12,13,14} Express each set in set builder notation.please help I don't understand these :(

OpenStudy (anonymous):

looks like integers between 7 and 14

OpenStudy (anonymous):

you could say \[\{x\in \mathbb{Z}| 7\leq x\leq 14\}\]

OpenStudy (anwaarullah):

x belong to R i guess

OpenStudy (anonymous):

where did you get R from?

OpenStudy (anwaarullah):

{ x member of reals | x ≥ 7 and x ≤ 14

OpenStudy (anwaarullah):

\[x \ge7 \le14\]

OpenStudy (zzr0ck3r):

There are an infinite amount of real numbers between 7 and 14, I would suggest what @satellite73 said.